OpenStudy (anonymous):

The chemical equation below shows the reaction between carbon dioxide (CO2) and lithium hydroxide (LiOH). CO2 + 2LiOH mc011-1.jpg Li2CO3 + H2O The molar mass CO2 is 44.01 g/mol. How many moles of LiOH are needed to react completely with 25.5 g of CO2? 0.290 0.579 1.16 1.73 @Somy @sweetburger @iPwnbunnies

OpenStudy (anonymous):

CO2 + 2LiOH = Li2CO3 + H2O

OpenStudy (anonymous):

isnt it 0.0290

OpenStudy (somy):

did u calculate it or guessed?

OpenStudy (anonymous):

guessed dont understand it

OpenStudy (somy):

not gonna work

OpenStudy (somy):

alright r u familiar with this formula mole= mass/Molecular mass?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

hhh well then im telling you mole= mass/Molecular mass both mass and molecular masses for CO2 are given put them in the formula and get mole for CO2

OpenStudy (anonymous):

so 1.16?

OpenStudy (somy):

~_~ how did u get it?

OpenStudy (anonymous):

i guesssed dont undestand this unit at all

OpenStudy (somy):

=_= hhh mole= 25.5/44.01 =0.579 mole of CO2 and ratio between CO2 and LiOH is 1:2 1----0.579mole 2---- x x= 2*0579 =

OpenStudy (anonymous):

x= 2*0579 = gives me 1.158

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

yes yes

OpenStudy (anonymous):

thats not an option tho

OpenStudy (somy):

well if you round it it'll be in the option ~_~

OpenStudy (anonymous):

1.16

OpenStudy (somy):

yup)